OpenStudy (anonymous):

The equation of the chord of the ellipse 2x2+5y2=20 which is bisected at the point (2,1) is

OpenStudy (anonymous):

Solve the stimulaneous equations \(y=mx+c \) \(2x^2+5y^2=20 \) Afte that, notice that it is bisected at (2,1).This should tell you that the mean of their x-coordinates (of the two points) =2. i.e \( \frac{x_1 + x_2}{2}=2\) This also tells you to use the sum of roots formula. Do you need further assistance?